# A Convex Hull Cheapest Insertion Heuristic for the Non-Euclidean TSP

### v3

@@ Abstract
The convex hull cheapest insertion heuristic is known to produce good solutions to the Traveling Salesperson Problem in Euclidean spaces, but it has not been extended to the non-Euclidean case. The proposed adaptation uses multidimensional scaling to first project the points into a Euclidean space, thereby enabling the generation of the convex hull that initializes the algorithm. To evaluate the proposed algorithm, non-Euclidean spaces are created by adding impassable separators to the TSPLIB benchmark data-set, or by using the $\mathcal{L}_{1}$ norm as a metric. This adapted heuristic is demonstrated to outperform the commonly used Nearest Neighbor heuristic and Nearest Insertion heuristic in 89% and 99% of the cases studied, respectively. When the genetic algorithm and ant colony optimization algorithms are provided 1 minute of computation time, the proposed heuristic tour costs are lower than the mean metaheuristic solutions found in 87% and 95% of the instances, respectively.
@@ 1 Introduction
The Traveling Salesperson Problem (TSP) involves finding the shortest possible tour that visits a set of locations exactly once before returning to the starting location [1]. In contrast to TSPs in Euclidean spaces, the non-Euclidean TSP includes obstacles in the environment, or a cost function that is not simply the pairwise Euclidean distance between locations [2]. For example, if the cost to be minimized is the time spent traveling between locations in a city, the optimal tour is dependent on the available road infrastructure that causes path deviations from the straight-line path between locations [3]. In a city with a rectangular grid of streets, the $\mathcal{L}_{1}$ norm, or the Manhattan norm, is sometimes used instead of the Euclidean distance [4].
When the number of locations to be visited is large, exact methods for computing the optimal solution to TSPs are intractable due to their $\mathcal{NP}$-hard nature [5, 6, 7]. When TSP solutions have to be found quickly, heuristics that rapidly find reasonably good solutions are typically used [8]. They are also used to provide solutions that act as upper bounds or as a warm start when initializing exact methods for faster convergence to the optimal solution [9, 10]. However, effective heuristics for the non-Euclidean TSP have been neglected in literature [11, 12], and the simple Nearest Neighbor (NN) or Nearest Insertion (NI) greedy heuristics are commonly used [13, 14].
@@
The Convex Hull Cheapest Insertion (CHCI) heuristic has been shown to produce superior solutions to greedy heuristic in most Euclidean test instances [15, 16]. The CHCI heuristic is initiated by a subtour created from the convex hull of locations, and its interior points are then progressively incorporated to the subtour in increasing order of insertion cost, until the complete tour is obtained. The initiation of the candidate subtour with the convex hull of the TSP points is advantageous because points on the boundary of the convex hull are visited in the same cyclic order as they appear in the optimal Euclidean TSP tour [17, 18]. However, the CHCI heuristic has not been adapted to the non-Euclidean TSP because it is initialized with the convex hull of points in a Euclidean space. While a common approach is to either neglect obstacles, or to replace the non-Euclidean cost function with Euclidean distances [19], this approximation is not always acceptable, especially when the true costs deviate significantly from the Euclidean distance [11]. In this context, using methods developed for Euclidean TSPs results in sub-optimal solutions [11].
@@
The contribution of this paper is the extension of the Euclidean CHCI algorithm to the non-Euclidean TSP, motivated by the expected reduction in tour cost when compared to the NN and NI heuristics. This is achieved by first applying multidimensional scaling (MDS) to find the set of points in a projected Euclidean space whose pairwise distances approximate the non-Euclidean pairwise cost. The points on the boundary of their convex hull are used to initiate a sub-tour onto which the remaining points are added in increasing order of their true non-Euclidean insertion cost ratios. While MDS has recently been applied to the non-Euclidean TSP for tour length estimation, local clustering, and to understand human cognition [20], the use of MDS to initiate the CHCI heuristic for the non-Euclidean TSP is a novel approach. After describing the algorithm and the heuristic and metaheuristic benchmarks, this paper reports the outcomes of extensive computational experiments on modified TSPLIB benchmark instances [21]
@@ 2 Non-Euclidean TSP
Consider a complete directed graph represented by $\mathcal{G}:=(V,A)$ where $V$ is the set of locations or nodes, and the directed arc set $A:=\{(v_{i},v_{j})|v_{i},v_{j}\in V,\forall i\neq j\}$ connects every ordered pair of distinct nodes in $V$. A cost function $c:A\rightarrow\mathbb{R}^{+}$ defines the metric to be minimized, and each arc $(v_{i},v_{j})\in A$ is associated with a defined cost $c(v_{i},v_{j})\in\mathbb{R}^{+}$. The objective of the TSP is to find the minimum cost sequence of consecutive arcs on the graph $\mathcal{G}$ that visits every node in $V$ exactly once and returns to the starting node. The resulting sequence is called a minimum cost Hamiltonian tour, and if $|V|=n$, it can be expressed as a sequence $T=(v_{1},v_{2},...,v_{n},v_{1})$ whose tour cost is the sum of costs of constituting arcs, given by $J=\textstyle\sum_{r=1}^{n}c(v_{r},v_{r+1})$. A constraint $v_{n+1}=v_{1}$ enforces that the starting and ending node are the same location.
@@ 3 Adapted Convex Hull Cheapest Insertion Algorithm
The adapted CHCI (ACHCI) algorithm first uses MDS to find a set of points in 2D space whose pairwise Euclidean distance approximates the non-Euclidean arc cost function $c$. This is initiated by assigning one of the TSP points to be the origin in a new Euclidean space, and then finding the coordinates of the remaining $n-1$ points such that their pairwise Euclidean distances are exactly their non-Euclidean costs [22, 23, 24, 25, 26]. Assuming a full rank cost matrix $C$, these points are first obtained in an $n-1$ dimension Euclidean space, after which they are projected to a 2D space.
Let the desired Euclidean coordinate equivalent of the node $v_{i}\in V$ be represented by $x_{i}\in\mathbb{R}^{n-1}$. In this paper, the notation $x_{i}\leftrightarrow v_{i}$ is used to indicate this mapping. For some $x_{j},x_{k}\in\mathbb{R}^{n-1}$, the relative position vectors of $x_{i}$ and $x_{j}$ with respect to $x_{k}$ are $(x_{i}-x_{k})$ and $(x_{j}-x_{k})$ respectively. If the Euclidean distance between these points is identical to their respective non-Euclidean costs as defined in cost matrix $C$, it can be verified that their inner product satsifies
@@
$\widetilde{X}=\widetilde{\Sigma}\widetilde{Q}^{T}$ (4)
@@
Figure 2: ACHCI subtour initiation and complete tour for the non-Euclidean TSP instance with 25 points
@@
For the example set of 25 non-Euclidean points, the subtour is initiated using the convex hull boundary points of $\widetilde{X}$ as shown in Fig. LABEL:Im1c:_CHullInit, after which the Euclidean coordinate approximation is discarded. The insertion cost ratios of Step 2, 3 and 4 use the true non-Euclidean cost matrix $C$ to obtain the complete tour, as shown in Fig. LABEL:Im1d:_ACHCIsol.
@@ 4 Benchmark algorithms
Two heuristic algorithms and two metaheuristic algorithms are chosen as benchmarks for comparison with the ACHCI algorithm. The NN and NI heuristic algorithms are based on fast and simple greedy selection rules [27] that are known to perform well [28] in constrained TSP applications [13, 14, 29, 30, 31]. Among metaheuristic algorithms that are commonly seen in TSP literature, the Genetic Algorithm (GA) and Ant Colony Optimization (ACO) are chosen for their ability to utilize and update a population of solutions, offering a broader exploration of the solution space [32].
@@ 4.2 Nearest Insertion Heuristic
    Find consecutive nodes $v_{i},v_{j}\in V$ in the subtour and $v_{k}\in V$ not in the subtour, that minimizes non-Euclidean insertion cost $C_{ik}+C_{kj}-C_{ij}$.
@@ 4.3 Genetic Algorithm
GA is a popular metaheuristic algorithm, inspired by the mechanisms of natural selection and genetics[33]. The GA process is initialized with a population of $m$ randomly generated tours, known as chromosomes. The quality of each chromosome is defined by its fitness function, assigned as the inverse of the TSP tour length. In subsequent iterations, these chromosomes are modified to find solutions with higher fitness value. At any iteration, the incumbent solution is defined as the chromosome with the highest fitness values, or lowest tour cost found up to that iteration. To ensure that the incumbent solution does not degrade as the iterations proceed, a defined number of high fitness chromosomes, called the elite population are carried over to the next generation without modification.
@@
The elite population along with the Roulette Wheel selected chromosomes form the $m$ parents for the next generation. Of these parents, a fraction $p_{1}$ is randomly selected for the Ordered Crossover operation, where segments of two parent chromosomes are exchanged to produce two offspring. To ensure that duplicate cities do not feature in the resulting offspring, each offspring is constructed by choosing a sub-sequence of one parent while preserving the relative order of remaining cities of the other [35]. Next, a fraction $p_{2}$ of the resulting population is randomly sampled for the process of Swap Mutation whereby two random locations within a chromosome are selected and swapped to form a new offspring [36]. Of the resulting population, a fraction $p_{3}$ is then sampled for the Inversion Mutation, where a subtour segment of a parent chromosome is randomly chosen and its order is reversed [34].
@@ 4.4 Ant Colony Optimization
ACO is another nature-inspired metaheuristic algorithm that mimics how a colony of ants finds the shortest paths to food sources based on distance and the strength of pheromone trails laid by individual ants [37]. While the GA uses stochastic processes to vary tour costs without explicitly evaluating edges in the arc set $A$, the ACO assigns pheromone strengths and heuristic values to the edges to inform exploration.
The ACO is initialized by computing local heuristic values $\eta_{ij}$ for each arc $(v_{i},v_{j})\in A$, assigned as the inverse of the cost $c(v_{i},v_{j})$ for $v_{i}\neq v_{j}$, and zero otherwise. The pheromone value $\tau_{ij}~{}\forall~{}(v_{i},v_{j})\in A$ is initially assigned a negligible value of $10^{-4}$. In subsequent iterations, $\tau_{ij}$ will be updated to reflect the effect of edge $(v_{i},v_{j})$ on tour costs.
@@
The influence of pheromone and heuristic information on tour edge selection is controlled by parameters $\alpha$ and $\beta$ respectively. Let $c_{a}$ represent the cost of the tour sampled by ant $a$ of colony $\mathcal{A}$, obtained by sampling locations based on Eq.(5). After tours have been constructed for the ant colony, pheromone levels $\tau_{ij}$ on each edge $(v_{i},v_{j})$ are updated before the next iteration:
@@ 5.1 Non-Euclidean Test Cases
The ACHCI solution to the ‘t70’ instance outfitted with the $\mathcal{L}_{1}$ norm is shown in Fig. LABEL:t70_L1, where it can be seen that paths are only permissible along the $x$ and $y$ directions. This deviation from the Euclidean straight line path is also seen in the case with 4 added separators, as shown in Fig. LABEL:Im_4a:_NETSP. The distortion of the Euclidean space caused by non-Euclidean cost functions is quantified in literature [3, 40] by the deviation factor (DF):
@@ 5.2 Performance Analysis
To analyze the performance of the ACHCI heuristic, extensive computational experiments were conducted in a Matlab R2022a environment on an Intel Xeon E5-2680 v4 CPU clocked at 2.4 GHz at the Ohio Super Computer [42]. The utility of the approximate algorithms compared in this paper lies in their ability to obtain solutions quickly, and for this reason, each algorithm is executed up to a time limit of 60 seconds. The hyperparameters that define the metaheuristic algorithms influence the convergence of incumbent solutions and their values are listed in Table 1. To account for the stochasticity involved in the GA and ACO algorithms, each algorithm is repeated 28 times for each TSPLIB case. The distribution of solutions found upon termination acts as a measure of their effectiveness.
The convergence of GA and ACO incumbent solutions along with the 95% confidence interval is shown in Fig. LABEL:Im:roC100convergence for the case where the $\mathcal{L}_{1}$ norm is applied to the TSPLIB instance ‘roC100’ with 100 points. The GA only uses tour cost as a metric for tour improvement which results in rapid rates of convergence. Contrarily, because the ACO updates pheromone values $\tau_{ij}$ of each edge, the computational demand per iteration is higher. While this results in a slower convergence rate, its incumbent solution costs are initially lower than that of GA. For an increased number of points, as seen in Fig. LABEL:Im:roB150convergence where 4 separators were added to the 150 points of instance ‘roB150’, the initial performance of the ACO is still superior to the GA.

### v5

@@ Abstract
The convex hull cheapest insertion heuristic produces good solutions to the Euclidean Traveling Salesperson Problem, but it has never been extended to the non-Euclidean problem. This paper uses multidimensional scaling to first project the points from a non-Euclidean space into a Euclidean space, enabling the generation of a convex hull that initializes the algorithm. To evaluate the proposed algorithm, non-Euclidean spaces are created by adding separators to the TSPLIB data-set, or by using the $\mathcal{L}_{1}$ norm as a metric.
@@ 1 Introduction
The Traveling Salesperson Problem (TSP) involves finding the shortest possible tour that visits a set of locations exactly once before returning to the starting location [1]. In contrast to TSPs in Euclidean spaces, the non-Euclidean TSP includes environmental obstacles, or a cost function that is not simply the pairwise Euclidean distance between locations [2]. For example, if the cost to be minimized is the time spent traveling between locations in a city, the optimal tour is dependent on the available road infrastructure that causes path deviations from the straight-line path between locations [3]. In case of a rectangular grid of streets, the $\mathcal{L}_{1}$ norm or the Manhattan norm is often used instead of the Euclidean distance [4, 5].
@@
When the size of the problem increases, i.e., for a large number of locations, exact methods for computing the optimal solution to TSPs are intractable due to their $\mathcal{NP}$-hard nature [6]. To accelerate exact methods by providing a solution as a warm start [7], or when TSP solutions have to be found in real time, heuristic algorithms are used to rapidly find reasonably good solutions [8]. Among heuristics developed for Euclidean problems, the Convex Hull Cheapest Insertion (CHCI) heuristic has been shown to produce superior solutions to greedy heuristics [9, 10]. The CHCI heuristic is initiated by a subtour of the convex hull of locations, and its interior points are then progressively incorporated into the subtour in increasing order of insertion cost, until the complete tour is obtained. The initiation of the candidate subtour with the convex hull of the TSP points is advantageous because points on the boundary of the convex hull are visited in the same cyclic order as they appear in the optimal Euclidean TSP tour [11, 12].
The initialization of the CHCI heuristic by the convex hull of points has prevented the utilization of this promising heuristic in non-Euclidean TSPs. In general, it is noted that effective heuristics for the non-Euclidean TSP have been neglected in literature, and the simple Nearest Neighbor (NN) or Nearest Insertion (NI) heuristics are commonly used instead [13]. Another approach is to either neglect obstacles, or to replace the non-Euclidean cost function with Euclidean distances [14]. However, this approximation is not always acceptable, especially when the true costs deviate significantly from the Euclidean distance [15]. In this context, using methods developed for Euclidean TSPs results in sub-optimal solutions [13].
@@
The contribution of this paper is the extension of the Euclidean CHCI algorithm to the non-Euclidean TSP, motivated by the expected reduction in tour cost when compared to the NN and NI heuristics. This is achieved by first applying multidimensional scaling (MDS) to find the set of points in a projected Euclidean space, whose pairwise distances approximate the non-Euclidean pairwise cost [16, 17]. The points on the boundary of their convex hull are used to initiate a sub-tour onto which the remaining points are added in increasing order of their true non-Euclidean insertion cost ratios. While MDS has recently been applied to the non-Euclidean TSP for tour length estimation, local clustering, and to understand human cognition [18, 19, 20], the use of MDS to initiate the CHCI heuristic for the non-Euclidean TSP is a novel approach.
After defining the proposed algorithm and its competing benchmark algorithms, this paper reports the outcomes of extensive computational experiments on TSPLIB benchmark instances that are modified by adding separators or using the $\mathcal{L}_{1}$ norm to set up non-Euclidean test cases [21]. This novel heuristic is demonstrated to outperform the NN and NI heuristic in 88% and 99% of the cases studied, respectively. When compared with metaheuristic algorithms, the proposed heuristic’s tour costs are lower than the solutions found by the genetic algorithm and ant colony optimization algorithm in 87% and 95% of the instances respectively.
@@ 2 Non-Euclidean TSP
Consider a complete directed graph represented by $\mathcal{G}:=(V,A)$ where $V$ is the set of locations or nodes, and the directed arc set $A:=\{(v_{i},v_{j})|v_{i},v_{j}\in V,\forall i\neq j\}$ which connects every ordered pair of distinct nodes in $V$. A cost function $c:A\rightarrow\mathbb{R}^{+}$ defines the metric to be minimized, and each arc $(v_{i},v_{j})\in A$ is associated with a defined cost $c(v_{i},v_{j})\in\mathbb{R}^{+}$. The objective of the TSP is to find the minimum cost sequence of arcs on the graph $\mathcal{G}$ that visits every node in $V$ exactly once and returns to the starting node. The resulting sequence is called a minimum cost Hamiltonian tour, and if $|V|=n$, it can be expressed as a sequence $T=(v_{1},v_{2},...,v_{n},v_{1})$ whose tour cost is the sum of costs of constituting arcs, given by $J=\textstyle\sum_{r=1}^{n}c(v_{r},v_{r+1})$. A constraint $v_{n+1}=v_{1}$ enforces that the starting and ending node are the same location.
@@ 3 Adapted Convex Hull Cheapest Insertion Algorithm
The proposed algorithm is referred to as the adapted CHCI (ACHCI) algorithm. It first uses MDS to find a set of points in 2D space whose pairwise Euclidean distance approximates the non-Euclidean arc cost function $c$. This is initiated by assigning one of the TSP points to be the origin in a new Euclidean space, and then finding the coordinates of the remaining $n-1$ points such that their pairwise Euclidean distances are exactly their non-Euclidean costs. Assuming a full rank cost matrix $C$, these points are first obtained in an $n-1$ dimension Euclidean space, after which they are projected to a 2D space.
Let the Euclidean coordinate equivalent of the node $v_{i}\in V$ be represented by $x_{i}\in\mathbb{R}^{n-1}$. In this paper, the notation $x_{i}\leftrightarrow v_{i}$ is used to indicate this mapping. For some $x_{j},x_{k}\in\mathbb{R}^{n-1}$, the relative position vectors of $x_{i}$ and $x_{j}$ with respect to $x_{k}$ are $(x_{i}-x_{k})$ and $(x_{j}-x_{k})$ respectively. If the Euclidean distance between these points is identical to their respective non-Euclidean costs as defined in cost matrix $C$, it can be verified that their inner product satsifies
@@
$\widetilde{X}=\widetilde{\Sigma}\widetilde{Q}^{T}$
@@
Figure 2: ACHCI subtour initiation and complete tour for a non-Euclidean TSP instance with 25 points
@@
For the example set of 25 non-Euclidean points, the subtour is initiated using the convex hull boundary points of $\widetilde{X}$ as shown in Fig. LABEL:Im1c:_CHullInit, after which the Euclidean coordinate approximation is discarded. The insertion cost ratios of Step 3, 4 and 5 use the true non-Euclidean cost matrix $C$ to obtain the complete tour, which is shown in Fig. LABEL:Im1d:_ACHCIsol.
@@ 4 Benchmark algorithms
Two heuristic algorithms and two metaheuristic algorithms are chosen as benchmarks for comparison with the ACHCI algorithm. The NN and NI heuristic algorithms are greedy algorithms based on selection rules and are known to perform well in constrained TSP applications[22, 23, 24]. Among metaheuristic algorithms that are commonly seen in TSP literature, the Genetic Algorithm (GA) and Ant Colony Optimization (ACO) are chosen as benchmarks because of their ability to utilize and update a population of solutions, offering a broader exploration of the solution space [25].
@@ 4.2 Nearest Insertion Heuristic
    Find consecutive nodes $(v_{i},v_{j})$ in the subtour, and $v_{k}$ not in the subtour, that minimizes non-Euclidean insertion cost $C_{ik}+C_{kj}-C_{ij}$.
@@ 4.3 Genetic Algorithm
The GA is a popular metaheuristic algorithm, inspired by the mechanisms of natural selection and genetics [26]. The GA process is initialized with a population of $m$ randomly generated tours, known as chromosomes. The quality of each chromosome is defined by its fitness function, assigned as the inverse of the TSP tour length. In subsequent iterations, these chromosomes are modified to find solutions with higher fitness value. At any iteration, the incumbent solution is defined as the chromosome with the highest fitness values or lowest tour cost found up to that iteration. To ensure that the incumbent solution does not degrade as the iterations proceed, a defined number of high fitness chromosomes, called the elite population are carried over to the next generation without modification.
@@
The elite population and the Roulette Wheel selected chromosomes form the $m$ parents for the next generation. Of these parents, a fraction $p_{1}$ is randomly selected for the Ordered Crossover operation, where segments of two parent chromosomes are exchanged to produce two offspring. To ensure that duplicate cities do not feature in the resulting offspring, each offspring is constructed by choosing a sub-sequence of one parent while preserving the relative order of remaining cities of the other [28]. Next, a fraction $p_{2}$ of the resulting population is randomly sampled for the process of Swap Mutation whereby two random locations within a chromosome are selected and swapped to form a new offspring [29]. Of the resulting population, a fraction $p_{3}$ is then sampled for the Inversion Mutation, where a subtour segment of a parent chromosome is randomly chosen and its order is reversed [27].
@@ 4.4 Ant Colony Optimization
The ACO is another nature-inspired metaheuristic algorithm that mimics how a colony of ants finds the shortest paths to food sources based on distance and the strength of pheromone trails laid by individual ants [30]. While the GA uses stochastic processes to vary tour costs without explicitly evaluating edges in the arc set $A$, the ACO assigns pheromone strengths and heuristic values to the edges to inform exploration.
The ACO is initialized by computing local heuristic values $\eta_{ij}$ for each arc $(v_{i},v_{j})\in A$, assigned as the inverse of the cost $c(v_{i},v_{j})$ for $v_{i}\neq v_{j}$, and zero otherwise. The pheromone value $\tau_{ij}~{}\forall~{}(v_{i},v_{j})\in A$ is initially assigned a negligible value called the primary tracing, and in subsequent iterations, $\tau_{ij}$ will be updated to reflect the effect of edge $(v_{i},v_{j})$ on tour costs.
@@
The influence of pheromone and heuristic information on tour edge selection is controlled by exponential parameters $\alpha$ and $\beta$ respectively. Let $c_{a}$ represent the cost of the tour sampled by ant $a$ of colony $\mathcal{A}$, obtained by sampling locations based on Eq.(4). After tours have been constructed for the ant colony, pheromone levels $\tau_{ij}$ on each edge $(v_{i},v_{j})$ are updated before the next iteration:
@@ 5.1 Non-Euclidean Test Cases
The ACHCI solution to the ‘t70’ instance outfitted with the $\mathcal{L}_{1}$ norm is shown in Fig. LABEL:t70_L1, where it can be seen that paths are only permissible along the $x$ and $y$ directions. A deviation from the Euclidean straight line path is also seen in the case with 4 added separators, as shown in Fig. LABEL:Im_4a:_NETSP. The distortion of the Euclidean space caused by non-Euclidean cost functions is quantified in literature [3, 32] by the deviation factor (DF):
@@ 5.2 Performance Analysis
To analyze the performance of the ACHCI heuristic, extensive computational experiments were conducted in a Matlab R2022a environment on an Intel Xeon E5-2680 v4 CPU clocked at 2.4 GHz at the Ohio Super Computer [34]. The utility of the approximate algorithms compared in this paper lies in their ability to obtain solutions quickly, and for this reason, each algorithm is executed up to a time limit of 60 seconds. The hyperparameters that define the metaheuristic algorithms influence the convergence of incumbent solutions. Their values are listed in Table 1. To account for the stochasticity involved in the GA and ACO algorithms, each algorithm is repeated 28 times for each TSPLIB case and the distribution of solutions is noted upon termination at 60 seconds.
The convergence of GA and ACO incumbent solutions along with their 95% confidence interval is shown in Fig. LABEL:Im:roC100convergence for the case where the $\mathcal{L}_{1}$ norm is applied to the TSPLIB instance ‘roC100’ with 100 points. The GA only uses tour cost as a metric for tour improvement which results in rapid rates of convergence. Contrarily, because the ACO samples locations and updates pheromone values $\tau_{ij}$ of each edge as per Eq. (4) and (5) respectively, the computational demand per iteration is higher. While this results in a slower convergence rate, its incumbent solution costs are initially lower than that of GA. In Fig. LABEL:Im:roB150convergence where 4 separators were added to the 150 points of instance ‘roB150’, it is seen that despite the larger number of points, the initial performance of the ACO is still superior.
Performance comparisons are conducted for 57 TSPLIB instances, with the number of points varying from 51 to 1,400. For each TSPLIB instance, one non-Euclidean case is created by assigning costs using the $\mathcal{L}_{1}$ norm, and 3 cases are created by adding 4, 16 or 64 impassable separators. For the sake of brevity, only 14 instances are tabulated in Table 2, though an unabridged version can be found in [35].